OpenStudy (anonymous):

Use DeMoivre's Theorem to find the fourth roots of: 16(cos(4pi/3) + isin(4pi/3))

OpenStudy (anonymous):

divide by 4

OpenStudy (anonymous):

well divide the angle by 4, also take the fourth root of 16, which is 2

OpenStudy (anonymous):

easy right?

OpenStudy (anonymous):

the angle

OpenStudy (anonymous):

which in this case is real real easy since \[\frac{4\pi}{3}\div 4=\frac{\pi}{3}\]

OpenStudy (anonymous):

after that you have a choice

OpenStudy (dan815):

This thm comes from eulers identity

OpenStudy (anonymous):

you can divide up the unit circle in two 4 equal parts, with one at \(\frac{\pi}{3}\) or you can keep adding \(2\pi\) and dividing first method is loads easier

OpenStudy (anonymous):



OpenStudy (anonymous):

ok i got it, but how would i use the thrm because my teacher marks points off if we dont what the directions says. @satellite73

OpenStudy (anonymous):

*if we dont do what the directions says

OpenStudy (anonymous):

can you help @dan815

OpenStudy (dan815):

yea what do u need

OpenStudy (anonymous):

i have to use the demoivres thrm, but idk how?

OpenStudy (dan815):

oky it comes from eulers identity

OpenStudy (anonymous):

idk what that is

OpenStudy (dan815):

okay u need some basic stuff first

OpenStudy (dan815):

There is a way to write a linear transformation matrix in terms of a multiplication of complex numbers

OpenStudy (dan815):

complex numbers are quite useful in this area to figure out rotations and scaling

OpenStudy (dan815):

so first thing u shud know is

OpenStudy (dan815):

if u have a vectors <1,1> & <1,2> lets say

OpenStudy (dan815):